# Disentangled Representation Learning with the Gromov-Monge Gap

### v2

@@ Abstract
Learning disentangled representations from unlabelled data is a fundamental challenge in machine learning. Solving it may unlock other problems, such as generalization, interpretability, or fairness. Although remarkably challenging to solve in theory, disentanglement is often achieved in practice through prior matching. Furthermore, recent works have shown that prior matching approaches can be enhanced by leveraging geometrical considerations, e.g., by learning representations that preserve geometric features of the data, such as distances or angles between points. However, matching the prior while preserving geometric features is challenging, as a mapping that fully preserves these features while aligning the data distribution with the prior does not exist in general. To address these challenges, we introduce a novel approach to disentangled representation learning based on quadratic optimal transport. We formulate the problem using Gromov-Monge maps that transport one distribution onto another with minimal distortion of predefined geometric features, preserving them as much as can be achieved. To compute such maps, we propose the Gromov-Monge-Gap (GMG), a regularizer quantifying whether a map moves a reference distribution with minimal geometry distortion. We demonstrate the effectiveness of our approach for disentanglement across four standard benchmarks, outperforming other methods leveraging geometric considerations.
@@ 1 Introduction
In this work, we demonstrate how to effectively combine geometric desiderata with prior matching within the VAE framework, using optimal transport (OT) theory (Santambrogio, 2015, Peyré and Cuturi, 2019). By treating mappings from the data manifold to the latent space (encoders) or vice versa (decoders) as transport maps $T:\mathcal{X}\to\mathcal{Y}$, we can leverage the Gromov-{Monge, Wasserstein} paradigm (Sturm, 2020, Mémoli, 2011), which aligns two distributions by finding a mapping that minimizes the distortion between intra-domain cost functions defined on their supports. Specifically, we consider cost functions $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})$ on $\mathcal{X}$ and $c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})$ on $\mathcal{Y}$ that encode geometric features such as scaled distances or angles. Consequently, the resulting mapping transforms one distribution onto the other while preserving these geometric features as much as possible.
@@
    Across four standard disentangled representation learning benchmarks, we show that incorporating geometry-preserving desiderata via the GMG significantly enhances disentanglement across various methods—from the standard $\beta$-VAE to the combination of $\beta$-TCVAE with HFS (Roth et al., 2023)—notably outperforming other methods that leverage geometric constraints.
2 Background and Related Works
@@ 2.1 Disentangled Representation Learning
Recent research suggests that disentanglement can arise by imposing geometric constraints on representation spaces (Horan et al., 2021). Most works (Karras et al., 2020, Huh et al., 2023) promote latent representations to preserve scaled distances between data points. Closest to our work, Lee et al. (2022) use a Jacobian-based regularization within a VAE, while Nakagawa et al. (2023) employ a distortion-based regularization, directly comparing pairwise squared Euclidean distances, within a Wasserstein AE (WAE, Tolstikhin et al. (2019)).
@@ 2.2 Quadratic Optimal Transport
OT (Peyré and Cuturi, 2019) involves transferring a probability distribution to another with inductive biases. When these distributions lie on incomparable domains, one relies on the Gromov-Monge (GM) and GW problems, also known as quadratic formulations of OT. Consider two compact $\mathcal{X}\subset\mathbb{R}^{d_{\mathcal{X}}}$, $\mathcal{Y}\subset\mathbb{R}^{d_{\mathcal{Y}}}$, each of them equipped with an intra-domain cost $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$. For technical reasons, we assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are CPD kernels (see Def. (A.1)). For $p\in\mathcal{P}(\mathcal{X})$ and $q\in\mathcal{P}(\mathcal{Y})$—two distributions supported on each domain—the GM problem (Mémoli and Needham, 2022) seeks a map $T:\mathcal{X}\to\mathcal{Y}$ that push-forwards $p$ onto $q$, while minimizing the distortion of the costs:
@@
When it exists, we call a solution $T^{\star}$ to (GMP) a Gromov-Monge map for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$. However, this formulation is ill-suited for discrete distributions $p,q$, as the constraint set might be empty in that case. Replacing maps by coupling $\pi\in\Pi(p,q)$, i.e. distributions on $\mathcal{X}\times\mathcal{Y}$ with marginals $p$ and $q$, we obtain the GW problem (Mémoli, 2011, Sturm, 2020)
@@
with $\mathbf{C}_{\mathcal{X}}=[c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})]_{i,i^{\prime}}$, $\mathbf{C}_{\mathcal{Y}}=[c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}})]_{j,j^{\prime}}\in\mathbb{R}^{n\times n}$, $U_{n}=\{\mathbf{P}\in\mathbb{R}^{n\times n}_{+},\mathbf{P}\mathbf{1}_{n}=\mathbf{P}^{T}\mathbf{1}_{n}=\tfrac{1}{n}\mathbf{1}_{n}\}$ and $H(\mathbf{P})=-\sum_{i,j=1}^{n}\mathbf{P}_{i,j}\log(\mathbf{P}_{i,j})$. As $\varepsilon\to 0$, we recover $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{0}=\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}$. In addition to yielding greater convexity (Rioux et al., 2023) and better statistical (Zhang et al., 2023) and properties, entropic regularization improves computational performance. We can solve (EGWP) using a scheme that iterates the Sinkhorn algorithm (Peyré et al., 2016). This solver has $\mathcal{O}(n^{2})$ memory complexity. Its time complexity depends on $c_{\mathcal{X}},c_{\mathcal{Y}}$. For general $c_{\mathcal{X}},c_{\mathcal{Y}}$, it runs in $\mathcal{O}(n^{3})$. However, for the most common practical choices of $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$ or $c_{\mathcal{X}}=c_{\mathcal{Y}}=\|\cdot-\cdot\|_{2}^{2}$, it can also be reduced to $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$, as detailed by Scetbon et al. (2022, §3 & Alg. 2).
Neural GW.
While for classical OT, numerous neural methods have been proposed (Makkuva et al., 2020, Korotin et al., 2022, Tong et al., 2023, Eyring et al., 2024), the GW setting has received less attention. Carrasco et al. (2024) propose an approach for solving GM with inner product costs, that involves a min-max-min optimization procedure. Klein et al. (2024) propose a framework to computeentropic GW couplings. Yet, these formulations are end-to-end, and not based on a regularizer that can be flexibly added to a loss function. Therefore, they cannot be combined with other models, as we do in this work with the $\beta$-VAE. In a concurrent work, Sotiropoulou and Alvarez-Melis (2024) introduced the Gromov-Monge gap. However, (i) they do not apply it for disentangled representational learning, and (ii) they do not investigate its properties, as we do in this work.
3 Disentanglement with the Gromov-Monge gap
@@
This section outlines our novel OT-based approach to disentanglement through geometric considerations within the VAE framework. First, we explore how an arbitrary map $T:\mathcal{X}\to\mathcal{Y}$ between two domains can be encouraged to preserve predefined geometric features. In a VAE, $T$ may represent either the encoder $e_{\phi}$, which generates latent codes from the data, or the decoder $d_{\theta}$, which reconstructs the data from these codes. In the case of the encoder, the source domain $\mathcal{X}$ corresponds to the data, and the target domain $\mathcal{Y}$ is the latent space, while for the decoder, these roles are swapped. Assuming that $d_{\theta}$ perfectly reconstructs the data from the latents produced by $e_{\phi}$, i.e., $e_{\phi}\circ d_{\theta}=\mathrm{I}d$, the preservation of geometric features by either $e_{\phi}$ or $d_{\theta}$ becomes equivalent. Therefore, in the following sections, we refer to $T$ as either the encoder or the decoder without loss of generality.
Outline.
Building on this perspective, this section begins by defining cost functions to encode geometric features and the notion of distortions in §3.1. We leverage this concept in §3.2 to introduce the Gromov-Monge Gap (GMG), a regularizer that measures whether a map moves distributions while preserving geometric features as much as possible, i.e., minimizing distortion. §3.3 then shows how the GMG can be efficiently computed from samples to be practically applicable in the VAE framework. This transitions into §3.4 studying (weak) convexity properties of the GMG, as an operator. Finally, in §3.5, we describe how to integrate the GMG with disentangled representation learning objectives, effectively combining prior matching with geometric constraints.
3.1 From the distortion…
@@
We encode the geometric features of interest using two cost functions, one for each domain: $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$. We aim for $T$ to preserve the geometric feature, which means that $T$ preserves these costs, that is, $c_{\mathcal{X}}(\mathbf{x},\mathbf{x^{\prime}})\approx c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x^{\prime}}))$ for $\mathbf{x},\mathbf{x^{\prime}}\in\mathcal{X}$. In practice, two types of cost function from this family are particularly relevant:
@@
  • [ii]
[…]
@@
$\mathrm{DST}_{r}(T)$ quantifies how much $T$ distorts geometric features encoded by $c_{\mathcal{X}},c_{\mathcal{Y}}$ on the support of $r$, that is, when $\mathrm{DST}_{r}(T)=0$, one has $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$ for $\mathbf{x},\mathbf{x}^{\prime}\in\mathrm{Spt}(r)$. Nakagawa et al. (2023) suggest promoting geometry preservation by regularizing the encoder $d_{\theta}$ using the DST, with $\mathbf{ScL2^{2}}$ as costs, and the latent representation as reference distribution, namely $r=e_{\phi}\sharp p_{\textrm{data}}$. While they use it within a WAE, translated in our VAE setting, this corresponds to minimizing
$\min_{\theta,\phi}\mathbb{E}_{\mathbf{x}\sim p_{\textrm{data}},\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}[\underbrace{-\log p_{\theta}(\mathbf{x}|\mathbf{z})}_{\textrm{(i) reconstruction}}+\underbrace{\beta D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))}_{\textrm{(ii) prior matching}}]+\underbrace{\lambda\,\mathrm{DST}_{r}(d_{\theta})}_{\textrm{(iii) geom. preservation}},$ (2)
where $\lambda>0$. Given the choice of the costs, $d_{\theta}$ is distortion-free (i.e., $\mathrm{DST}_{r}(d_{\theta})=0$), if it is a scaled isometry. Since a scaled isometry that maps the prior onto the data distribution may not exist, there is an inherent trade-off between minimizing terms (ii) and (iii), which are responsible for achieving practical disentanglement. As these terms cannot be simultaneously minimized to 00, the DST loss will move away from accurately matching the prior, which will negatively impact the quality of the learned latent representations. This naturally raises the question of how to avoid this over-penalization. Instead of seeking a distortion-free decoder, we should seek a decoder that transports the prior to the data distribution with minimal distortion of the costs $c_{\mathcal{X}},c_{\mathcal{Y}}$. In other words, the decoder should be a Gromov-Monge map between the prior and the data distributions for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$ (see 2.2). Conversely, if we choose to regularize the encoder, the same reasoning applies by swapping the roles of the prior and the data distribution. In the next section, we introduce a regularizer to fit Gromov-Monge maps, which we will use as a replacement to the DST.
3.2 …to The Gromov-Monge Gap

### v3

@@ Abstract
Learning disentangled representations from unlabelled data is a fundamental challenge in machine learning. Solving it may unlock other problems, such as generalization, interpretability, or fairness. Although remarkably challenging to solve in theory, disentanglement is often achieved in practice through prior matching. Furthermore, recent works have shown that prior matching approaches can be enhanced by leveraging geometrical considerations, e.g., by learning representations that preserve geometric features of the data, such as distances or angles between points. However, matching the prior while preserving geometric features is challenging, as a mapping that fully preserves these features while aligning the data distribution with the prior does not exist in general. To address these challenges, we introduce a novel approach to disentangled representation learning based on quadratic optimal transport. We formulate the problem using Gromov-Monge maps that transport one distribution onto another with minimal distortion of predefined geometric features, preserving them as much as can be achieved. To compute such maps, we propose the Gromov-Monge-Gap (GMG), a regularizer quantifying whether a map moves a reference distribution with minimal geometry distortion. We demonstrate the effectiveness of our approach for disentanglement across four standard benchmarks, outperforming other methods leveraging geometric considerations. Code is available at: https://github.com/ExplainableML/GMG
@@ 1 Introduction
In this work, we demonstrate how to effectively combine geometric desiderata with prior matching within the VAE framework, using optimal transport (OT) theory (Santambrogio, 2015, Peyré and Cuturi, 2019). By treating mappings from the data manifold to the latent space (encoders) or vice versa (decoders) as transport maps $T:\mathcal{X}\to\mathcal{Y}$, we can leverage the Gromov-{Monge, Wasserstein} paradigm (Sturm, 2023, Mémoli, 2011), which aligns two distributions by finding a mapping that minimizes the distortion between intra-domain cost functions defined on their supports. Specifically, we consider cost functions $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})$ on $\mathcal{X}$ and $c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})$ on $\mathcal{Y}$ that encode geometric features such as scaled distances or angles. Consequently, the resulting mapping transforms one distribution onto the other while preserving these geometric features as much as possible.
@@
    Across four standard disentangled representation learning benchmarks, we show that incorporating geometry-preserving desiderata via the GMG significantly enhances disentanglement across various methods, from the standard $\beta$-VAE to the combination of $\beta$-TCVAE with HFS (Roth et al., 2023).
2 Background: On Disentanglement, Quadratic-OT and Distortion
@@ 2.1 Disentangled Representation Learning
Recent studies have revealed a fundamental connection between geometric structure preservation and disentanglement in learned representations (Gropp et al., 2020, Chen et al., 2020a, Lee et al., 2022, Nakagawa et al., 2023, Huh et al., 2023). This connection was theoretically established by Horan et al. (2021) proving that unsupervised disentanglement is always feasible when the generative factors are sufficiently non-Gaussian and maintain local isometry to the data. Our work builds directly on this insight by developing a learning framework that promotes representations that are as close as possible to being isometric to the data. To quantify geometric preservation between spaces of different dimensions, we leverage quadratic OT theory, which originated in Koopmans and Beckmann (1957) and was formalized by Mémoli (2011) as a framework for measuring isometric correspondence between metric spaces. We detail these tools in Section 2.2 before showing how they can be used to learn representations in Section 2.3. Additionally, in a concurrent work, Sotiropoulou and Alvarez-Melis (2024) also introduced the Gromov-Monge gap. However, they do not apply it for disentangled representational learning.
@@ 2.2 Quadratic Optimal Transport
OT (Peyré and Cuturi, 2019) theory studies efficient ways to map a probability distribution onto another. Linear OT formulations, such as the Monge (1781) problem, require domains $\mathcal{X},\mathcal{Y}$ that can be directly compared through a cost function $c(x,y)$ defined between their elements. When these distributions lie on incomparable domains, one must instead rely on quadratic formulations of OT (Q-OT), which instead compare geometric structure through intra-domain costs, also known as the Gromov-Monge (GM) and GW problems. In the context of representation learning, representation and data spaces are incomparable by design, which necessitates the use of Q-OT in this work.
@@
Consider two compact $\mathcal{X}\subset\mathbb{R}^{d_{\mathcal{X}}}$, $\mathcal{Y}\subset\mathbb{R}^{d_{\mathcal{Y}}}$, equipped with intra-domain cost $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$. We assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$) are CPD kernels (Def. (A.1)). For $p\in\mathcal{P}(\mathcal{X})$ and $q\in\mathcal{P}(\mathcal{Y})$, two distributions supported on each domain, the GM problem (Mémoli and Needham, 2022) seeks a map $T:\mathcal{X}\to\mathcal{Y}$ that push-forwards $p$ onto $q$, while minimizing the distortion of the costs:
@@
When it exists, we call a solution $T^{\star}$ to (GMP) a Gromov-Monge map for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$. However, this formulation is ill-suited for discrete distributions $p,q$, as the constraint set might be empty in that case. Replacing maps by coupling $\pi\in\Pi(p,q)$, i.e. distributions on $\mathcal{X}\times\mathcal{Y}$ with marginals $p$ and $q$, we obtain the GW problem (Mémoli, 2011, Sturm, 2023)
@@
with $\mathbf{C}_{\mathcal{X}}=[c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})]_{i,i^{\prime}}$, $\mathbf{C}_{\mathcal{Y}}=[c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}})]_{j,j^{\prime}}\in\mathbb{R}^{n\times n}$, $U_{n}=\{\mathbf{P}\in\mathbb{R}^{n\times n}_{+},\mathbf{P}\mathbf{1}_{n}=\mathbf{P}^{T}\mathbf{1}_{n}=\tfrac{1}{n}\mathbf{1}_{n}\}$ and $H(\mathbf{P})=-\sum_{i,j=1}^{n}\mathbf{P}_{i,j}\log(\mathbf{P}_{i,j})$. As $\varepsilon\to 0$, we recover $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{0}=\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}$. Entropic regularization improves computational performance, as we can solve (EGWP) using a scheme that iterates the Sinkhorn algorithm (see Appendix D.3 for full details). This solver has $\mathcal{O}(n^{2})$ memory complexity. Its time complexity depends on $c_{\mathcal{X}},c_{\mathcal{Y}}$. For general $c_{\mathcal{X}},c_{\mathcal{Y}}$, it runs in $\mathcal{O}(n^{3})$. However, for the most common practical choices of $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$ or $c_{\mathcal{X}}=c_{\mathcal{Y}}=\|\cdot-\cdot\|_{2}^{2}$, it can also be reduced to $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$, as detailed by Scetbon et al. (2022, §3 & Alg. 2).
2.3 Distortion in Representation Learning
@@
Given an arbitrary map $T:\mathcal{X}\to\mathcal{Y}$, we consider how it can be learned to preserve predefined geometric features. In a VAE, $T$ may represent either the encoder $e_{\phi}$, which generates latent codes from the data, or the decoder $d_{\theta}$, which reconstructs the data from these codes. In the case of the encoder, $\mathcal{X}$ corresponds to the data, and $\mathcal{Y}$ is the latent space, while for the decoder, these roles are swapped. Assuming that $d_{\theta}$ perfectly reconstructs the data from the latents produced by $e_{\phi}$, i.e., $e_{\phi}\circ d_{\theta}=\mathrm{I}d$, the preservation of geometric features by either $e_{\phi}$ or $d_{\theta}$ becomes equivalent. Therefore, in the following sections, we refer to $T$ as either the encoder or the decoder without loss of generality.
@@
We encode geometric features using a cost function for each domain: $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$. Ideally, $T$ should preserve geometry, which means that $T$ preserves costs, that is, $c_{\mathcal{X}}(\mathbf{x},\mathbf{x^{\prime}})\approx c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x^{\prime}}))$ for $\mathbf{x},\mathbf{x^{\prime}}\in\mathcal{X}$. In practice, two types of costs are often used:
@@
  • [$i$]
    Cosine-Similarity: $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=\textrm{cos-sim}(\mathbf{x},\mathbf{x}^{\prime})\vcentcolon=\langle\tfrac{\mathbf{x}}{\|\mathbf{x}\|_{2}},\tfrac{\mathbf{x}^{\prime}}{\|\mathbf{x}^{\prime}\|_{2}}\rangle$ and $c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})=\textrm{cos-sim}(\mathbf{y},\mathbf{y}^{\prime})$ similarly. On has $\textrm{cos-sim}(\mathbf{x},\mathbf{x}^{\prime})=\cos(\theta_{\mathbf{x},\mathbf{x}^{\prime}})$ where $\theta_{\mathbf{x},\mathbf{x}^{\prime}}$ is the angle between $\mathbf{x}$ and $\mathbf{x}^{\prime}$. A map $T$ preserving $c_{\mathcal{X}},c_{\mathcal{Y}}$ then preserves the angles between the points, i.e. it is a conformal map. Note that if $T$ is (scaled) isometry (see above), it is a conformal map.
@@
$\mathrm{DST}_{r}(T)$ quantifies how much $T$ distorts geometric features encoded by $c_{\mathcal{X}},c_{\mathcal{Y}}$ on the support of $r$, that is, when $\mathrm{DST}_{r}(T)=0$, one has $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})=c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$ for $\mathbf{x},\mathbf{x}^{\prime}\in\mathrm{Spt}(r)$.
Distortion as a Loss for Representation Learning. Nakagawa et al. (2023) suggest promoting geometry preservation by regularizing the encoder $d_{\theta}$ using the DST, with $\mathbf{ScL2^{2}}$ as costs, and the latent representation as reference distribution, namely $r=e_{\phi}\sharp p_{\textrm{data}}$. While they use it within a WAE, we translate their objective to VAE setting adopted later in the paper. This results in:
$\min_{\theta,\phi}\mathbb{E}_{\mathbf{x}\sim p_{\textrm{data}},\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}[\underbrace{-\log p_{\theta}(\mathbf{x}|\mathbf{z})}_{\textrm{(i) reconstruction}}+\underbrace{\beta D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))}_{\textrm{(ii) prior matching}}]+\underbrace{\lambda\,\mathrm{DST}_{r}(d_{\theta})}_{\textrm{(iii) geom. preservation}},\quad\lambda>0$ (2)
Given the choice of the costs, $d_{\theta}$ is distortion-free (i.e., $\mathrm{DST}_{r}(d_{\theta})=0$), if it is a scaled isometry.
Challenges Arising from a Mixed Loss. Since a scaled isometry that maps the prior onto the data distribution may not exist, there is an inherent trade-off between minimizing terms (ii) and (iii), which are responsible for achieving practical disentanglement. As these terms cannot be simultaneously minimized to 0, the DST loss will move away from accurately matching the prior, which will negatively impact the quality of the learned latent representations. This naturally raises the question of how to avoid this over-penalization. Instead of seeking a distortion-free decoder, we should seek a decoder that transports the prior to the data distribution with minimal distortion of the costs $c_{\mathcal{X}},c_{\mathcal{Y}}$. In other words, the decoder should be a Gromov-Monge map between the prior and the data distributions for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$ (see 2.2). Conversely, if we choose to regularize the encoder, the same reasoning applies by swapping the roles of the prior and the data distribution. In the next section, we introduce a regularizer to fit Gromov-Monge maps, which we will use as a replacement to the DST.
3 Disentanglement with the Gromov-Monge gap
Building on these geometric preservation principles, we introduce in §3.1 the Gromov-Monge Gap (GMG), a regularizer that measures whether a map moves distributions while preserving geometric features as much as possible, i.e., minimizing distortion while fitting the marginal constraints. §3.2 then shows how the GMG can be efficiently computed from samples to be practically applicable in the VAE framework. This transitions into §3.3 studying (weak) convexity properties of the GMG, as an operator. Finally, in §3.4, we describe how to integrate the GMG with disentangled representation learning objectives, effectively combining prior matching with geometric constraints.
3.1 An Efficient Gap Formulation for Distortion